OpenStudy (enricoanders):

Find the equation in standard form of the line with slope -5/4 that passes through the point (5,4) Hint: Remember that A, B, and C must be integers, not fractions

OpenStudy (enricoanders):

@agent0smith

OpenStudy (enricoanders):

can someone please explain this to me please

OpenStudy (shaima):

what do u think it is

OpenStudy (enricoanders):

I dont know, im truly lost

OpenStudy (enricoanders):

5x-5/4y=4

OpenStudy (shaima):

okay ill help

OpenStudy (shaima):

alright so what do u think we have to do.

OpenStudy (mhchen):

Solution: 4y+5x =41 Steps: I used point-slope form (y-4) = (-5/4)(x-5) Substitution 4y-16 = -5x+25 Multiplication and Distribution Property of Equality 4y+5x = 41 Addition Property of Equality You can check with https://www.desmos.com/calculator